# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

### v2

@@ 5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to assess our explainability method across different text generation models and different scenarios.
5.2.1 Accuracy
@@
To evaluate the performance of instruction-based text generation models, we tested ten different prompts and their variations through 64 perturbations. Heatmaps were generated to display the weights of each keyword in the prompts across various models, including OpenAI GPT, LLAMA, and Claude-AI. A ground truth was defined, assigning weights of 1 to critical keywords in the prompts and 0 to less relevant or auxiliary words. These ground truth values were compared with the extracted heatmaps to assess the interpretability and accuracy of each model.
@@
To quantify accuracy, we employed multiple metrics, including Attention Accuracy (ATT ACC), F1-Score for Attention (ATT F1), and Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC). The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
Among the evaluated models, Claude-AI demonstrated the strongest performance, achieving the highest scores in all three metrics: 0.82 in ATT ACC, 0.67 in ATT F1, and 0.88 in ATT AUROC. These results indicate Claude-AI’s superior ability to consistently focus on essential input tokens across varying prompts. In contrast, LLAMA achieved the second-highest accuracy but showed a notably lower F1-score (0.40), suggesting potential inconsistency in identifying key tokens. OpenAI GPT yielded balanced results but lagged slightly behind Claude-AI in all metrics, highlighting room for improvement in its attention fidelity. Ll, Claude-AI provides more reliable and semantically focused attention compared to its counterparts.
Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
@@
models, evaluated under two conditions: different prompts and different images.
Model name ATT ACC ATT F1 ATT AUROC
OpenAI-GPT 0.70 0.59 0.84
LLaMA 0.76 0.40 0.84
Claude-AI 0.82 0.67 0.88
5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added only the token  at the end of the text, extracted the coefficients again, and computed the Jaccard index between the coefficients before and after the perturbation. The averaged Jaccard indices across all ten prompts for each model are reported in Table 2. Higher scores indicate greater stability under minor input changes.
@@ 5.2.2 Stability
Table 2: Stability scores across models based on average Jaccard index over ten prompts
@@
OpenAI-GPT 0.62
LLaMA 0.45
Claude-AI 0.44
5.2.3 Consistency
To compute consistency, we ran the same code with 64 perturbations for each prompt over 10 iterations, ensuring that the outputs remained consistent across these repetitions for text generation models such as OpenAI GPT, LLAMA, and Claude-AI, as shown in Table 3. We computed variance and standard deviation metrics for each word coefficient. This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
Among all models, OpenAI-GPT exhibited the highest consistency, achieving a variance of 0.0000 and a standard deviation of 0.0046. This indicates that its predictions are the most stable across repeated runs, confirming its robustness and reliability in generating interpretable outputs.
Table 3: Consistency metrics for different models on the prompt
“What is the meaning of life?”
@@ 5.2.3 Consistency
OpenAI-GPT 0.0000 0.0046
LLaMA 0.0048 0.0663
Claude-AI 0.0011 0.0331
5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
[…]
Table 4: Fidelity metrics for different models on the prompt
“What is the meaning of life?”
[…]
OpenAI-GPT 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
LLaMA 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
From the results, we observe that Claude-AI achieves the lowest WMSE, WMAE, and mean-L1, indicating superior performance in minimizing error. On the other hand, OpenAI-GPT performs best in terms of $R^{2}_{\omega}$, mean-L2, and adjusted $R^{2}_{\hat{\omega}}$, demonstrating stronger alignment and generalization of the surrogate model. LLaMA ranks in the middle across most metrics but does not lead in any. This suggests that different models excel in different fidelity dimensions, highlighting the importance of multi-metric evaluation when interpreting black-box model behavior.
5.2.5 Fidelity Across Different number of Text Perturbations
Tables 56, and 7 present the results of fidelity computations for varying numbers of perturbations across the OpenAI GPT, LLAMA, and Claude-AI models, respectively. These results are derived using a weighted linear regression as the surrogate model and the Wasserstein distance for computing distances. The evaluation metrics include weighted mean squared error (WMSE), weighted mean absolute error (WMAE), coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$), and average loss metrics such as mean-L1 and mean-L2.
Table 5: Performance metrics for different numbers of perturbations in OpenAI GPT
@@ 5.2.5 Fidelity Across Different number of Text Perturbations
Table 6: Performance metrics for different numbers of perturbations in LLaMA
@@
Table 7: Performance metrics for different numbers of perturbations in Claude-AI
@@
For OpenAI GPT and LLaMA, increasing the number of perturbations generally improves fidelity across most metrics, with 256 perturbations achieving the best WMSE, WMAE, and loss values. Interestingly, for Claude-AI, 64 perturbations yield the lowest WMSE and WMAE, while 256 achieves the best loss (mean-L2 and mean-L1). However, the highest $R^{2}$ scores for all three models occur at 32 perturbations, suggesting that beyond a certain point, more perturbations may dilute interpretability despite improving local fit. Therefore, there is a trade-off between the number of perturbations and the reliability of the surrogate model, depending on the metric of interest.
5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
Fidelity is also computed by comparing various distance metrics for text generation models such as OpenAI GPT using text-to-text comparisons. The surrogate models used are Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME), as shown in Table 8. We explore different combinations of distance metrics, including Cosine similarity and Wasserstein Distance (WD). Fidelity is measured using WMSE, WMAE, mean-$L_{1}$, mean-$L_{2}$, and coefficient of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$).
Table 8: Fidelity results for different distance measures with 30 perturbations
WLR Fidelity Metrics
@@ 5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
Cosine WD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
WD WD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
WD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
Our Proposed Method: gSMILE.
Based on the results in Table 8, the combination where Wasserstein Distance (WD) is used for both input similarity and output shift consistently yields the highest values of the weighted coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$). Specifically, underweighted linear regression, the WD–WD configuration achieves $R^{2}_{\omega}=\textbf{0.7104}$ and $R^{2}_{\hat{\omega}}=\textbf{0.6409}$, which significantly outperform other configurations.
Motivated by these findings, we introduce our method, called gSMILE (generalized SMILE), which utilizes the Wasserstein distance for both input and output comparisons. This choice enhances the surrogate model’s fidelity and stability in capturing local shifts in model behavior.
We use gSMILE throughout subsequent sections as the foundation for explainability evaluations on large language models (LLMs). These include assessments of fidelity, consistency, and stability, ensuring the proposed approach is both theoretically grounded and empirically adequate.
@@ 5.2.7 Computation Complexity
We evaluate the computational efficiency of our proposed method, gSMILE, using three instruction-based text generation models: OpenAI GPT, LLaMA, and Claude-AI. Although these models differ in scale and architecture, the objective here is not to compare the models themselves but rather to assess the performance of different explainability methods applied to each model under identical hardware and runtime conditions. All evaluations are conducted with 60 perturbations per input prompt.
We compare three explainability approaches—LIME, gSMILE, and Bay-LIME—based on their execution time. As reported in Table 9, gSMILE consistently requires the highest runtime across all model frameworks. This is mainly due to its reliance on the Wasserstein distance to compute similarity between output distributions, which involves solving high-dimensional optimization problems with time complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [145]. By contrast, LIME and Bay-LIME utilize cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
Despite the higher computational cost, gSMILE delivers significantly better fidelity, stability, and robustness to adversarial perturbations—essential properties for trustworthy explanations. These advantages are supported by results in Tables 8 and 2 and align with findings from prior work on SMILE [11], which demonstrated superior robustness against adversarial input variations.
These findings suggest that gSMILE’s computational overhead is justified by the quality of explanations it provides, making it highly suitable for applications where interpretability, consistency, and resistance to manipulation are paramount.
Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
@@
OpenAI-GPT 156.82 170.70 161.43
LLaMA 113.00 118.99 116.40
Claude-AI 355.09 372.95 349.88
@@ 5.2.8 Token-Level Attention for Different Sentence Structures
To investigate how sentence structure affects token-level attention in large language models (LLMs), we analyzed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts. These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
For each sentence, we extracted the attention values directly from the final layer of the LLM. We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualized using a box plot (see Figure 12), where each point represents a unique sentence. Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.

### v4

@@ 5.2 Quantitative Results
We quantitatively evaluate the proposed explainability method using multiple metrics, ATT Accuracy, ATT Stability, ATT Consistency, ATT Fidelity, and computational complexity, across different text generation models and scenarios. These metrics provide a multi-faceted assessment, capturing not only the accuracy of attributions but also their robustness, reproducibility, and efficiency in practical applications.
5.2.1 Attribution Accuracy
@@
We assessed attribution accuracy by testing ten distinct prompts and their variations, each subjected to 64 perturbations. For each prompt, we generated heatmaps showing the weight assigned to every keyword across three models: OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1. A ground-truth attribution map was defined by assigning a weight of 1 to critical keywords and 0 to auxiliary or less relevant tokens. Model-generated heatmaps were then compared to this ground truth.
@@
Accuracy was quantified using three complementary metrics: Attention Accuracy (ATT ACC), Attention F1-Score (ATT F1), and the Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC). Scores were averaged over all prompts, and the results are reported in Table 1.
Anthropic’s Claude 2.1 achieved the highest scores across all metrics, 0.82 ATT ACC, 0.67 ATT F1, and 0.88 ATT AUROC, indicating a strong and consistent focus on the most relevant tokens. Meta’s LLaMA 3.1 Instruct Turbo (70B) ranked second in accuracy but had a notably lower ATT F1 (0.40), suggesting occasional inconsistency in identifying all key tokens within a prompt. OpenAI’s gpt-3.5-turbo-instruct showed balanced but slightly lower scores compared to Claude 2.1.
These findings highlight that while all models can identify necessary tokens to some extent, Claude 2.1 demonstrates a more semantically precise and reliable attribution pattern, an essential property for trust in explainability-driven applications.
@@
Table 1: Performance metrics (ATT ACC, ATT F1, ATT AUROC, and published accuracy) for various models, evaluated under two conditions: different prompts and different images.
Model name ACC ATT ACC ATT F1 ATT AUROC
OpenAI’s gpt-3.5-turbo-instruct 0.70 0.70 0.59 0.84
Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.69 0.76 0.40 0.84
Anthropic’s Claude 2.1 0.79 0.82 0.67 0.88
5.2.2 Attribution Faithfulness
We assessed attribution faithfulness across the three evaluated large language models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, by measuring how well attribution scores align with the models’ externally reported factual accuracies. Table 1 presents both the published accuracies and our attribution metrics (ATT ACC, ATT F1, ATT AUROC).
The analysis reveals a clear positive relationship: models with higher factual accuracy tend to exhibit more faithful attributions. For instance, Claude 2.1, with a reported factual accuracy of 78.5%, achieves the highest ATT AUROC (0.88), reflecting strong alignment between its most essential tokens and its correct predictions. In Comparison, OpenAI’s gpt-3.5-turbo-instruct (70% accuracy) and LLaMA 3.1 (68.8% accuracy) both yield a slightly lower ATT AUROC of 0.84.
We quantified this trend by computing the Pearson correlation between published accuracy and each attribution metric (Table 2). ATT AUROC showed the most substantial alignment ($r\approx 0.994$), followed by ATT ACC ($r\approx 0.804$) and ATT F1 ($r\approx 0.802$).
These results indicate that attribution faithfulness can serve as a practical proxy for overall model reliability. When a model performs better on external benchmarks, it is also more likely to assign importance to semantically relevant input tokens. This makes ATT AUROC, in particular, a valuable diagnostic for selecting models in high-stakes interpretability scenarios.
Table 2: Pearson correlation coefficients ($r$) between published accuracy and attribution metrics.
Metric Pearson $r$
ATT ACC 0.804
ATT F1 0.802
ATT AUROC 0.994
5.2.3 Attribution Stability
We measure attribution stability using the Jaccard index, which quantifies the overlap between sets of necessary tokens before and after a small, controlled input perturbation. For each of the three evaluated models, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, we selected ten prompts and recorded the token-level attribution scores. We then introduced a minimal change by appending the token  to the end of each prompt, recalculated the attributions, and computed the Jaccard index between the pre- and post-perturbation sets of essential tokens.
Table 3 reports the average stability scores across all prompts. Higher values indicate greater resilience of attribution patterns to irrelevant input changes. OpenAI’s gpt-3.5-turbo-instruct achieved the highest stability score (0.62), suggesting that its attention to key tokens is less likely to shift under trivial modifications. In contrast, both LLaMA 3.1 (0.45) and Claude 2.1 (0.44) exhibited lower stability, indicating greater sensitivity to small, semantically insignificant changes.
From an interpretability standpoint, higher ATT stability is desirable in real-world deployments: it means that explanations are less likely to fluctuate unpredictably when prompts are rephrased or slightly altered, improving trust and consistency in model behaviour.
@@ 5.2.3 Attribution Stability
Table 3: ATT stability scores across models based on average Jaccard index over ten prompts.
@@
OpenAI’s gpt-3.5-turbo-instruct 0.62
Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.45
Anthropic’s Claude 2.1 0.44
5.2.4 Attribution Consistency
Attribution consistency measures how reproducible token-level importance scores are when the same analysis is repeated under identical conditions. For each model, OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1, we ran 64 perturbations for a given prompt across 10 independent iterations. For each token, we computed the variance and standard deviation of its attribution score across runs. Low variance indicates that the model’s attribution patterns are stable and not influenced by randomness in the perturbation process or internal model initialisation.
Table 4 summarises the results using the example prompt: “What is the meaning of life?” OpenAI’s gpt-3.5-turbo-instruct achieved near-perfect reproducibility, with a variance of 0.0000 and a standard deviation of only 0.0046. In contrast, LLaMA 3.1 and Claude 2.1 displayed higher variability, with LLaMA showing the least consistency overall.
From an interpretability perspective, high ATT consistency is crucial: if attribution maps fluctuate unpredictably between identical runs, users cannot reliably trust the explanations. The results here suggest that, at least for this prompt, GPT-3.5 provides the most deterministic and reliable attributions.
Table 4: ATT consistency metrics for different models on the prompt “What is the meaning of life?”
@@ 5.2.4 Attribution Consistency
OpenAI’s gpt-3.5-turbo-instruct 0.0000 0.0046
Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.0048 0.0663
Anthropic’s Claude 2.1 0.0011 0.0331
5.2.5 ATT Fidelity for Different Instruction-Based Models
ATT fidelity measures how well a surrogate model can approximate the attribution patterns of a black-box model. For the prompt “What is the meaning of life?”, we evaluated OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 using 32 perturbations. Fidelity was computed with a weighted linear regression surrogate and Wasserstein distance, assessing WMSE, weighted $R^{2}$ ($R^{2}_{\omega}$), WMAE, mean $L_{1}$, mean $L_{2}$, and weighted adjusted $R^{2}$ ($R^{2}_{\hat{\omega}}$).
Table 5: ATT fidelity metrics for different models on the prompt “What is the meaning of life?”
Model name WMSE $R^{2}_{\omega}$ WMAE $L_{1} mean$-$L_{2}$ $R^{2}_{\hat{\omega}}$
OpenAI’s gpt-3.5-turbo-instruct 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
Meta’s LLaMA 3.1 Instruct Turbo (70B) 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
Anthropic’s Claude 2.1 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
The results reveal a trade-off between error minimisation and alignment quality. Anthropic’s Claude 2.1 achieves the lowest WMSE, WMAE, and mean-$L_{1}$, suggesting it produces the most faithful low-error surrogate approximations. Conversely, OpenAI’s gpt-3.5-turbo-instruct leads in $R^{2}_{\omega}$, mean-$L_{2}$, and adjusted $R^{2}_{\hat{\omega}}$, indicating stronger overall correlation and generalisation of the surrogate model to the original attribution patterns. LLaMA 3.1 shows mid-range performance across most metrics, without topping any category.
From a practical perspective, these findings underscore that fidelity is multi-dimensional: a model optimising for minimal error may not consistently achieve the highest correlation with its surrogate, and vice versa. Thus, a multi-metric approach is essential when selecting models for high-stakes interpretability tasks.
5.2.6 Attribution Fidelity Across Different Numbers of Text Perturbations
Tables 68 report ATT fidelity for OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1 using varying numbers of perturbations. Fidelity was measured with a weighted linear regression surrogate and Wasserstein distance, using WMSE, WMAE, $R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$, mean-$L_{1}$, and mean-$L_{2}$.
For gpt-3.5-turbo-instruct and LLaMA 3.1, larger perturbation counts (especially 256) generally reduced error metrics (WMSE, WMAE, $L_{1}, mean$-$L_{2}$), suggesting a finer-grained local approximation of model attributions. In contrast, Claude 2.1 achieved its lowest WMSE and WMAE at only 64 perturbations, while still benefiting from higher counts in some loss metrics.
Interestingly, all three models reached their highest $R^{2}_{\omega}$ at 32 perturbations, indicating that correlation-based alignment peaks early, while additional perturbations may improve local fit but dilute global attribution patterns.
From a practical standpoint, these results suggest a trade-off: more perturbations can yield more precise local surrogate fits but may reduce overall attribution alignment. For interpretability workflows, an intermediate number of perturbations (e.g., 64–128) may offer a balanced choice, depending on whether correlation or error minimisation is the primary goal.
Table 6: Performance metrics for different numbers of perturbations in OpenAI’s gpt-3.5-turbo-instruct
@@ 5.2.6 Attribution Fidelity Across Different Numbers of Text Perturbations
Table 7: Performance metrics for different numbers of perturbations in Meta’s LLaMA 3.1 Instruct Turbo (70B)
@@
Table 8: Performance metrics for different numbers of perturbations in Anthropic’s Claude 2.1
@@
5.2.7 Attribution Fidelity for Different Distance Metrics and Surrogate Models
We compared ATT fidelity across combinations of distance metrics and surrogate models for OpenAI’s gpt-3.5-turbo-instruct, using 30 perturbations. Surrogates included Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME) [70], and distance metrics included Cosine similarity, Wasserstein Distance (WMD), and their combination (WMD+C). Fidelity was evaluated using WMSE, WMAE, $L_{1}, mean$-$L_{2}$, and coefficients of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$), as shown in Table 9.
Table 9: ATT fidelity results for different distance measures with 30 perturbations.
WLR ATT Fidelity Metrics
@@ 5.2.7 Attribution Fidelity for Different Distance Metrics and Surrogate Models
Cosine WMD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
WMD WMD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
WMD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
WMD+C WMD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
Key Findings and the gSMILE Approach.
Across both surrogate models, using Wasserstein Distance for both input similarity and output shift consistently produced the most substantial alignment between the surrogate and the target model. In WLR, the WMD–WMD configuration achieved the highest $R^{2}_{\omega}$ (0.7104) and $R^{2}_{\hat{\omega}}$ (0.6409), substantially outperforming cosine-based approaches.
This result motivated our proposed method, gSMILE (generalised SMILE), which standardises on WMD–WMD for attribution fidelity. The Wasserstein Distance captures fine-grained distributional differences in token attributions, making it particularly effective for text-to-text comparisons.
By adopting WMD–WMD, gSMILE improves not only ATT fidelity but also stability across perturbations, providing a consistent basis for all subsequent evaluations (fidelity, consistency, stability) in this work. This design choice balances theoretical robustness with empirical performance, making it well-suited for LLM interpretability workflows.
@@ 5.2.8 Computation Complexity
We assessed the computational efficiency of gSMILE using three instruction-based text generation models: OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo (70B), and Anthropic’s Claude 2.1. The goal here is not to compare these models, but to evaluate how different explainability methods perform under identical hardware and runtime conditions, with 60 perturbations per input prompt.
We compared three approaches, LIME, gSMILE, and Bay-LIME, based on execution time. As shown in Table 10, gSMILE consistently requires the most runtime. This is primarily due to its use of the Wasserstein distance for comparing output distributions, which involves solving high-dimensional optimisation problems with complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [138]. By contrast, LIME and Bay-LIME rely on cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
Although gSMILE incurs a higher computational cost, it delivers substantially better ATT fidelity, ATT stability, and robustness to adversarial perturbations, critical for producing trustworthy explanations. These benefits are supported by the results in Tables 9 and 3 and are consistent with prior findings on SMILE [13], which showed improved robustness against adversarial input variations.
In applications where interpretability, stability, and resistance to manipulation are priorities, gSMILE’s additional runtime is a justified trade-off for higher-quality, more reliable explanations.
Table 10: Execution times (in seconds) for each framework and explainability method with 60 perturbations.
@@
OpenAI’s gpt-3.5-turbo-instruct 156.82 170.70 161.43
Meta’s LLaMA 3.1 Instruct Turbo (70B) 113.00 118.99 116.40
Anthropic’s Claude 2.1 355.09 372.95 349.88
@@ 5.2.9 Token-Level Attention for Different Sentence Structures
To investigate how sentence structure affects token-level attention in large language models (LLMs), we analysed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts. These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
For each sentence, we extracted the attention values directly from the final layer of the LLM. We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualised using a box plot (see Figure 13), where each point represents a unique sentence. Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.